TC=250+60Q-3Q2+0.10Q3 W-where TC is total cost and Q is quantity of output produced.

a)Write an algebraic expressions for the Average Total Cost ?
b)Determine the output level where AVC costs are minimized ?
c)Determine the output level where marginal cost are minimized ?
1
Expert's answer
2017-05-03T09:02:09-0400
a)ATC=250/Q+60-3Q+0.1

b)=-3+0.2Q

-3+0.2Q=0

Q=7.46

c)MC=60-6Q+0.2

-6+0.4Q

-6+0.4Q=0

Q=15